# Very Basic Set theory

### v6

@@ 10. Natural numbers
In (Halmos [1974] p.47) the Peano Axioms for the set of all natural numbers ”used to be considered as the fountainhead of all mathematical knowledge”. In this section we define ‘number sequences’ satisfying the Peano Axioms for natural numbers.
@@
$s$ is ‘transitive’ if $\forall u(u\in s\rightarrow u\subseteq s)$, or equivalently $\bigcup s\subseteq s$. $\Box$
@@
It is straightforward to define a number sequence satisfying the Peano Axioms for natural numbers. The first axiom of the Peano Axioms reads ‘there is a first natural number’ (denoted in (Mendelson [2015] p.153) by ‘00’).
@@
Any pair of individuals is a first number of a natural number sequence. Therefore ‘the set of all natural numbers’ is nonexisting. However, to postulate ‘set $\omega$ of all natural numbers with first number $\alpha$’ is not excluded. Then $\bigcup\omega=\omega$. Therefore we can define ‘transfinite ordinal numbers with first number $\omega$’ analogous to natural numbers (the finite ordinal numbers) with first number $\alpha$.
In the Introduction (Wittgenstein [1999] p.21) Bertrand Russell writes
”No logic can be considered adequate until it has been shown to be capable of dealing with transfinite numbers”.
The logic in Very basic set theory is capable to deal with transfinite numbers.
@@ Conclusion
Very basic set theory is a set theory that has its philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus. It is a set theory with individuals, the logic is capable to deal with transfinite numbers and the naming of individuals is logically first. But individuals are not needed for mathematical purposes therefore individuals are banished from ‘mainstream set theory’.

### v7

@@ 7. Natural numbers
In (Halmos [1974] p.47) the Peano Axioms for the set of all natural numbers ”used to be considered as the fountainhead of all mathematical knowledge”. In this section we define ‘transitive sets’ and ‘number sequences’ satisfying the Peano Axioms for natural numbers.
@@
$s$ is transitive if $\forall u(u\in s\rightarrow u\subseteq s)$, or equivalently $\bigcup s\subseteq s$. $\Box$
@@
It is straightforward to define a number sequence satisfying the Peano Axioms for natural numbers. The first axiom of the Peano Axioms reads ‘there is a first natural number’. Natural numbers are also known as ‘finite ordinals’.
@@
Any pair of individuals is a first number of a natural number sequence. Therefore ‘the set of all natural numbers’ is nonexisting. However, it is not excluded to postulate ‘set $\omega$ of all natural numbers with first number $\alpha$’. Then $\bigcup\omega=\omega$. Therefore we can define ‘transfinite ordinal numbers with first number $\omega$’ analogous to natural numbers with first number $\alpha$. Therefore the logic in Very basic set theory is capable to deal with transfinite numbers.
@@ Conclusion
Very basic set theory is a set theory that has its philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus. It is a set theory with individuals, the logic is capable to deal with transfinite numbers and the naming of individuals is logically first. There is no ‘empty set’ and regularity is not an axiom but a theorem.